# Assessing Omitted Variable Bias

### v5

@@ 3.3 Adding Assumptions on the Outcome Equation
This is the ratio of the standard deviation of the index of unobservables in the outcome equation (1) to the standard deviation of the index of observables in the same equation. Thus $r_{Y}$ measures the relative impact of unobservables and observables on outcomes, whereas $r_{X}$ measures their relative impact on treatment. We now consider a bound on this ratio $r_{Y}$.
@@
In our next result, which uses restrictions on both $r_{X}$ and $r_{Y}$, we assume that the omitted variable of interest $W_{2}$ is a scalar, and then normalize its variance to 1. We leave the extension to vector $W_{2}$ to future work. Let $\mathcal{B}_{I}(\bar{r}_{X},\bar{r}_{Y})$ denote the identified set for $\beta_{\text{long}}$ under this scalar $W_{2}$ assumption, A1, A3, and A5. Unlike the identified set of Theorem 2, this set is less analytically tractable. We characterize it in Appendix LABEL:sec:generalIdentSet. In this section we show that it is nonetheless very easy to numerically implement a breakdown analysis, which is often researchers’ main focus.
@@
Masten and Poirier (2020) call this the robust region because the conclusion of interest, $\beta_{\text{long}}\geq\underline{b}$, holds for any combination of sensitivity parameters in this region. The size of this region is therefore a measure of the robustness of our baseline conclusion.
[…]
@@
We can now state our second main result.
@@
    subject to $\text{subject to }\ p(z,\operatorname*{cov}(W_{1},c_{1}Y+c_{2}X);\bar{r}_{X})\geq 0$
@@
Theorem 3 shows that the breakdown frontier can be computed as the solution to a smooth optimization problem. Importantly, this problem only requires searching over a 4-dimensional space. In particular, this dimension does not depend on the dimension of the covariates $W_{1}$. Consequently, it remains computationally feasible even with a large number of observed covariates, as is often the case in empirical practice. For example, the results for our empirical application take about 15 seconds to compute.
@@
Recommendations. To assess the impact of different choices of $(\bar{r}_{X},\bar{r}_{Y})$ on the robustness of their results, researchers can plot sample analog estimates of the function $\bar{r}_{Y}^{\text{bf}}(\cdot,\underline{b})$ (for example, see the solid line in the right plots of figures 1 and 2 in section 5 below). The conclusion of interest $\beta_{\text{long}}\geq\underline{b}$ holds for all pairs $(\bar{r}_{X},\bar{r}_{Y})$ below this function. In practice it is also useful to report scalar measures of robustness to omitted variables. One way to do this while also incorporating restrictions on the impact of the omitted variables on outcomes is to set $\bar{r}_{X}=\bar{r}_{Y}$. This assumption says that the impact of unobservables relative to observables is the same for outcomes as it is for treatment. Under this common maximal impact assumption, the breakdown point for the common value of the sensitivity parameter is
@@ 3.4 Interpreting and Calibrating the Sensitivity Parameters
where $W=(W_{0},W_{1},W_{2})$. Likewise, write equation (2) as
@@
for $k=1,\ldots,d_{1}$, where $\pi_{1,\text{med}}=(\pi_{1,\text{med},k},\pi_{1,\text{med},-k})$ is the coefficient on the observed covariates $W_{1}=(W_{1k},W_{1,-k})$ from OLS of treatment $X$ on $(1,W_{1})$. $\rho_{k}$ measures the importance of variable $k$ relative to all other observed covariates. Moreover, unlike $r_{X}$, the $\rho_{k}$ values are point identified from the data. We suggest that researchers compare their estimates of the breakdown point $\bar{r}_{X}^{\text{bp}}$ against estimates of the values $\rho_{1},\ldots,\rho_{d_{1}}$. This approach to the calibration of sensitivity parameters was first proposed in Hsu and Small (2013) and is similar to suggestions in Masten and Poirier (2018) and Cinelli and Hazlett (2020). We illustrate this approach in our empirical analysis in section 5. Finally, here we defined $\rho_{k}$ by comparing a single covariate against all others. More generally one could compare the impact of a group of covariates against all others; we omit this extension for brevity.
@@ 3.5 Restricting the Controls’ Endogeneity
In some applications, the bounds obtained using restrictions only on $\bar{r}_{X}$ and $\bar{r}_{Y}$ only may be quite large, even for small values of these sensitivity parameters. In this case, researchers may be willing to restrict the relationship between the observed covariates and the omitted variable. So next we present identification results that allow researchers to impose these restrictions. To motivate these results, note that for scalar $W_{2}$ with $\operatorname*{var}(W_{2})=1$ we can write the omitted variable bias as a function of the coefficient $\gamma_{2}$ on $W_{2}$ in the long regression outcome equation (1) and the coefficient $\pi_{2}$ on $W_{2}$ in the selection equation (2) as follows:
@@
$R_{W_{2}\sim W_{1}}\in[\underline{c},\bar{c}]$ for known values of $[\underline{c},\bar{c}]\subseteq[0,1]$.
Our results in section 3 can be thought of as the no-restrictions case $[\underline{c},\bar{c}]=[0,1]$. We first consider restrictions under assumption A6 and a bound on $r_{X}$. We then consider restrictions on $r_{Y}$ as well. We conclude this section by discussing an approach to assessing control endogeneity and to calibrating the values $\underline{c}$ and $\bar{c}$.
@@
Recall the definition of $\bar{z}_{X}(\bar{r}_{X},\underline{c},\bar{c})$ from equation (LABEL:eq:barz). Note that $\bar{z}_{X}(\bar{r}_{X},\underline{c},\bar{c})=\bar{z}_{X}(\bar{r}_{X},0,1)$ when $\bar{r}_{X}\in[\underline{c},\bar{c}]$. As before, the sensitivity parameters will only affect the bounds via this function.
@@
The interpretation of Theorem 4 is similar to our earlier result Theorem 2. It characterizes the largest and smallest possible values of $\beta_{\text{long}}$ when some selection on unobservables is allowed and the controls are allowed to be partially but not arbitrarily endogenous. We also make no restrictions on the coefficients in the outcome equation. As before, with the exception of at most three singletons, the interval $[\underline{B}(\bar{r}_{X},\underline{c},\bar{c}),\overline{B}(\bar{r}_{X},\underline{c},\bar{c})]$ is the identified set for $\beta_{\text{long}}$ under these assumptions.
Earlier we saw that $\bar{r}_{X}<1$ is necessary for the bounds of Theorem 2 to be finite. Theorem 4 shows that, if we are willing to make assumptions on $R^{2}_{W_{2}\sim W_{1}}$ through $[\underline{c},\bar{c}]$, then we can allow for $\bar{r}_{X}>1$ while still obtaining finite bounds since $\bar{z}_{X}(\bar{r}_{X},\underline{c},\bar{c})$ is finite whenever $\bar{r}_{X}<\bar{c}^{-1}$. Thus there is a trade-off between (i) the magnitude of selection on unobservables we can allow for and (ii) the magnitude of control endogeneity. We can use breakdown frontiers to summarize these trade-offs. Specifically, when $\beta_{\text{med}}\geq 0$, define
@@
    subject to $\text{subject to }\ p(z,c_{1}\operatorname*{cov}(W_{1},Y)+c_{2}\operatorname*{cov}(W_{1},X);\bar{r}_{X})\geq 0$
@@
Theorem 3 is the special case obtained by setting $[\underline{c},\bar{c}]=[0,1]$.
@@ 4 Causal Models
In this section we describe four different causal models in which the parameter $\beta_{\text{long}}$ in equation (1) has a causal interpretation. These models are based on three different kinds of identification strategies: Unconfoundedness, difference-in-differences, and instrumental variables. Here we focus on simple models, but our analysis can be used anytime the causal parameter of interest can be written as the coefficient on a treatment variable in a long regression of the form in equation (1).
@@ 4.1 Unconfoundedness for Linear Models
Linear latent unconfoundedness can be interpreted linear parametric version of the nonparametric latent unconfoundedness assumption
@@ 4.2 Nonparametric Unconfoundedness
In the previous section we considered a linear parametric version of the unconfoundedness identification strategy. Here we consider a nonparametric version, which allows us to extend our analysis to models with heterogeneous treatment effects. Let $Y(x)$ denote potential outcomes, $X$ be a binary treatment, and $W=(W_{1},W_{2})$ a vector of covariates. Suppose the following two assumptions hold:
@@ 4.3 Difference-in-Differences
Let $Y_{t}(x_{t})$ denote potential outcomes at time $t$, where $x_{t}$ is a logically possible treatment value. For simplicity we do not consider models with dynamic effects of treatment or of covariates. Also suppose $W_{2t}$ is a scalar for simplicity. Suppose
@@
Consider the linear potential outcome model (7). Suppose the following exogeneity assumption holds:
  • $\operatorname*{cov}(\Delta X,\Delta V)=0$, $\operatorname*{cov}(\Delta W_{2},\Delta V)=0$, and $\operatorname*{cov}(\Delta W_{1},\Delta V)=0$.
Then $\beta_{c}=\beta_{\text{long}}$.
@@ 4.4 Instrumental Variables
where $U$ is an unobserved scalar random variable and $(\beta_{c},\gamma_{1},\gamma_{2})$ are unknown constants. Thus $\beta_{c}$ is the causal effect of $Z$ on $Y$. In an instrumental variables analysis, this is typically called the reduced form causal effect, and $Y(z)$ are reduced form potential outcomes. Suppose $\operatorname*{cov}(Z^{\perp W_{1},W_{2}},Y(z)^{\perp W_{1},W_{2}})=0$, a linear parametric version of conditional independence between the instrument $Z$ and the unobserved $U$ given $(W_{1},W_{2})$. Then $\beta_{c}$ equals the OLS coefficient on $Z$ from the long regression of $Y$ on $(1,Z,W_{1},W_{2})$. In this model, Theorem 1 implies that $\beta_{c}$ is also obtained as the coefficient on $Z$ in the medium regression of $Y$ on $(1,Z,W_{1})$, and thus is point identified. In this case, assumption A2 is an instrument exogeneity assumption. Our results in section 3 thus allow us to analyze the impacts of instrument exogeneity failure on conclusions about the reduced form causal effect of $Z$ on $Y$, $\beta_{c}$.
@@ 5.1 Data
Next we describe the observed covariates. We partition these covariates into $W_{1}$ and $W_{0}$ by following the implementation of Oster’s (2019) approach in Bazzi et al. (2020). $W_{1}$, the calibration covariates which are used to calibrate selection on unobservables, is a set of geographic and climate controls: Centroid Latitude, Centroid Longitude, Land area, Average rainfall, Average temperature, Elevation, Average potential agricultural yield, and Distance from the centroid to rivers, lakes, and the coast. $W_{0}$, the control covariates which are not used to calibrate selection on unobservables, includes state fixed effects. The questionnaire based outcomes use individual level data. For those analyses, we also include age, age-squared, gender, race, and survey wave fixed effects in $W_{0}$. In Bazzi et al. (2020), they were included in $W_{1}$. We instead include them in $W_{0}$ to keep the set of calibration covariates $W_{1}$ constant across the five main specifications. This allows us to directly compare the robustness of our baseline results across different specifications. We discuss the results that include state fixed effects in $W_{1}$ in appendix LABEL:sec:additionalEmpirical.
@@ 5.2 Baseline Model Results
[talltblr]rowsep=0pt {talltblr} [ caption = The Effect of Frontier Life on Opposition to Government Intervention and Redistribution., remarkNote = Panel A and the first row of Panel B replicate columns 1, 2, 4, 6, and 7 of table 3 in Bazzi, Fiszbein, and Gebresilasse (2020), while the second row of Panel B and Panel C are new. As in Bazzi et al. (2020), Panel B uses Oster’s rule of thumb choice $R_{Y\sim X,W_{1},W_{2}\mathrel{{\bullet}}W_{0}}^{2}=1.3\cdot\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$., ] p0.25¿ *3¿p0.15 — *2¿p0.15 Prefers Cut Public Spending on Poor Prefers Cut Public Spending on Welfare Prefers Reduce Debt by Spending Cuts County Property Tax Rate Republican Presidential Vote Share
@@
$\widehat{\bar{r}}_{X}^{\text{ bp}}$ (×$\times 100$) 2.81 3.05 5.85 72.6 80.4
$\widehat{\bar{r}}^{\text{ bp}}$ (×$\times 100$) 54.1 74.4 83.9 91.8 95.9

### v6

@@ 3.3 Adding Assumptions on the Outcome Equation
This is the ratio of the standard deviation of the index of unobservables in the outcome equation (1) to the standard deviation of the index of observables in the same equation. Thus $r_{Y}$ measures the relative impact of unobservables and observables on outcomes, whereas $r_{X}$ measures their relative impact on treatment. Note that, like the discussion of $r_{X}$ in section 3.1, $r_{Y}$ can also be interpreted as a ratio of causal effects under a linearity assumption on potential outcomes. We now consider a bound on this ratio $r_{Y}$.
@@
In our next result, which uses restrictions on both $r_{X}$ and $r_{Y}$, we assume that the omitted variable of interest $W_{2}$ is a scalar, and then normalize its variance to 1. We leave the extension to vector $W_{2}$ to future work. Let $\mathcal{B}_{I}(\bar{r}_{X},\bar{r}_{Y})$ denote the identified set for $\beta_{\text{long}}$ under this scalar $W_{2}$ assumption, A1, A3, and A5. Unlike the identified set of Theorem 1, this set is less analytically tractable. We characterize it in Appendix C. In this section we show that it is nonetheless very easy to numerically implement a breakdown analysis, which is often researchers’ main focus.
@@
This set is called the robust region because the conclusion of interest, $\beta_{\text{long}}\geq\underline{b}$, holds for any combination of sensitivity parameters in this region. The size of this region is therefore a measure of the robustness of our baseline conclusion.
Although we do not have a closed form expression for the smallest and largest elements of $\mathcal{B}_{I}(\bar{r}_{X},\bar{r}_{Y})$, our next main result shows that we can still easily compute the breakdown frontier numerically. To state the result, we first define some additional notation. Let $\|\cdot\|$ denote the Euclidean norm. Define the set
@@
For simplicity, we also normalize the treatment variable so that $\operatorname*{var}(X)=1$ and the covariates so that $\operatorname*{var}(W_{1})=I$. All of the results below can be rewritten without these normalizations at the cost of additional notation. We can now state our second main result.
@@
    $\text{subject to }\ p(z,\operatorname*{cov}(W_{1},c_{1}Y+c_{2}X);\bar{r}_{X})\geq 0$
@@
Theorem 2 shows that the breakdown frontier can be computed as the solution to a smooth optimization problem. Importantly, this problem only requires searching over a 4-dimensional space. In particular, this dimension does not depend on the dimension of the covariates $W_{1}$. Consequently, it remains computationally feasible even with a large number of observed covariates, as is often the case in empirical practice. For example, the results for our empirical application take about 15 seconds to compute on a standard machine.
@@
Practical Recommendations for Applied Researchers
To assess the impact of different choices of $(\bar{r}_{X},\bar{r}_{Y})$ on the robustness of their results, researchers can plot sample analog estimates of the function $\bar{r}_{Y}^{\text{bf}}(\cdot,\underline{b})$ (for example, see the solid line in the right plots of figures 1 and 2 in section 5 below). The conclusion of interest $\beta_{\text{long}}\geq\underline{b}$ holds for all pairs $(\bar{r}_{X},\bar{r}_{Y})$ below this function. In practice it is also useful to report scalar measures of robustness to omitted variables. One way to do this while also incorporating restrictions on the impact of the omitted variables on outcomes is to set $\bar{r}_{X}=\bar{r}_{Y}$. This assumption says that the maximal impact of unobservables relative to observables is the same for outcomes as it is for treatment. Under this common maximal impact assumption, the breakdown point for the common value of the sensitivity parameter is
@@ 3.4 Interpreting and Calibrating the Sensitivity Parameters
where $W\coloneqq(W_{0},W_{1},W_{2})$. Likewise, write equation (2) as
@@
for $k=1,\ldots,d_{1}$, where $\pi_{1,\text{med}}=(\pi_{1,\text{med},k},\pi_{1,\text{med},-k})$ is the coefficient on the observed covariates $W_{1}=(W_{1k},W_{1,-k})$ from OLS of treatment $X$ on $(1,W_{1})$. $\rho_{k}$ measures the importance of variable $k$ relative to all other observed covariates. Moreover, unlike $r_{X}$, the $\rho_{k}$ values are point identified from the data. We suggest that researchers compare their estimates of the breakdown point $\bar{r}_{X}^{\text{bp}}$ against estimates of the values $\rho_{1},\ldots,\rho_{d_{1}}$. We illustrate this approach in our empirical analysis in section 5. Similar approaches to empirically calibrating sensitivity parameters were previously proposed in Hosman et al. (2010), Hsu and Small (2013), Masten and Poirier (2018), and Cinelli and Hazlett (2020). Importantly, as Hosman et al. (2010) discuss in detail, values like $\rho_{k}$ are “reference points for speculation” about the true, but unknown and completely unidentified value of $r_{X}$. Put differently: There is no logical guarantee that $r_{X}$ will be smaller than any of the $\rho_{k}$’s. Instead, their purpose is simply to help us use data to reason about plausible values of the true, but unknown sensitivity parameter $r_{X}$. Finally, here we defined $\rho_{k}$ by comparing a single covariate against all others. More generally one could compare the impact of a group of covariates against all others; we omit this extension for brevity.
@@ 3.5 Restricting the Controls’ Endogeneity
Our results thus far have allowed for arbitrarily endogenous controls. In this section we present more general identification results that allow researchers to impose restrictions on the correlation between the observed covariates and the omitted variable. There are at least two reasons why researchers may be interested in imposing such restrictions: (1) First, it allows us to compare bounds obtained under exogenous controls with those obtained under arbitrarily endogenous controls. This allows us to assess the impact of endogenous controls on the robustness of one’s results. (2) Second, when the observed controls are believed to be sufficiently endogenous, it may be plausible to impose a lower bound on their correlation.
To motivate our formal results, note that for scalar $W_{2}$ with $\operatorname*{var}(W_{2})=1$ we can write the omitted variable bias as a function of the coefficient $\gamma_{2}$ on $W_{2}$ in the long regression outcome equation (1) and the coefficient $\pi_{2}$ on $W_{2}$ in the selection equation (2) as follows:
@@
$R_{W_{2}\sim W_{1}}\in[\underline{c},\bar{c}]$ for known values of $\underline{c},\bar{c}]\subseteq[0,1]$.
Our results in section 3 can be thought of as the agnostic case $\underline{c},\bar{c}]=[0,1]$. We first consider restrictions under Assumption A6 and a bound on $r_{X}$. We then consider restrictions on $r_{Y}$ as well. We conclude this section by discussing an approach to assessing control endogeneity and to calibrating the values $\underline{c}$ and $\bar{c}$.
@@
Appendix equation (LABEL:eq:z_X_formula) defines the function $\bar{z}_{X}(\bar{r}_{X},\underline{c},\bar{c})$; its exact form is unimportant here. The sensitivity parameters will only affect the bounds via this function.
@@
Theorem 1 can be obtained as a special case of Theorem 3 by setting $\underline{c}=0$ and $\overline{c}=1$, and hence the two theorems have similar interpretations: Theorem 3 characterizes the largest and smallest possible values of $\beta_{\text{long}}$ when some selection on unobservables is allowed and the controls are allowed to be partially but not arbitrarily endogenous. We also make no restrictions on the coefficients in the outcome equation. As before, with the exception of at most three singletons, the interval $\underline{B}(\bar{r}_{X},\underline{c},\bar{c}),\overline{B}(\bar{r}_{X},\underline{c},\bar{c})]$ is the identified set for $\beta_{\text{long}}$ under these assumptions.
Earlier we saw that $\bar{r}_{X}<1$ is necessary for the bounds of Theorem 1 to be finite. Theorem 3 shows that, if we are willing to make assumptions on $R^{2}_{W_{2}\sim W_{1}}$ through $\underline{c},\bar{c}]$, then we can allow for $\bar{r}_{X}>1$ while still obtaining finite bounds since $\bar{z}_{X}(\bar{r}_{X},\underline{c},\bar{c})$ is finite whenever $\bar{r}_{X}<\bar{c}^{-1}$. Thus there is a trade-off between (i) the magnitude of selection on unobservables we can allow for and (ii) the magnitude of control endogeneity. We can use breakdown frontiers to summarize these trade-offs. Specifically, when $\beta_{\text{med}}\geq 0$, define
@@
    $\text{subject to }\ p(z,c_{1}\operatorname*{cov}(W_{1},Y)+c_{2}\operatorname*{cov}(W_{1},X);\bar{r}_{X})\geq 0$
@@
Theorem 2 is the special case obtained by setting $\underline{c},\bar{c}]=[0,1]$.
@@ 4 Causal Models
While the results in section 3 are stated under general conditions, empirical researchers are often interested in assessing sensitivity to omitted variables in causal models. In this section, we review how $\beta_{\text{long}}$ in equation (1) can be interpreted in the context of four specific causal models. These models are based on three identification strategies: Unconfoundedness, difference-in-differences, and instrumental variables. Here we focus on simple models, but our analysis can be used anytime the causal parameter of interest can be written as the coefficient on a treatment variable in a long regression of the form in equation (1).
@@ 4.1 Unconfoundedness for Linear Models
Linear latent unconfoundedness can be interpreted as a linear parametric version of the nonparametric latent unconfoundedness assumption
@@ 4.2 Nonparametric Unconfoundedness
In the previous section we considered a linear parametric version of the unconfoundedness identification strategy. Here we consider a nonparametric version, which allows us to extend our analysis to models with heterogeneous treatment effects. Let $Y(x)$ denote potential outcomes, $X$ be a binary treatment, and $W\coloneqq(W_{1},W_{2})$ a vector of covariates. Suppose the following two assumptions hold:
@@ 4.3 Difference-in-Differences
Let $Y_{t}(x_{t})$ denote potential outcomes at time $t$, where $x_{t}$ is a logically possible treatment value. For simplicity we do not consider models with dynamic effects. Also suppose $W_{2t}$ is a scalar for simplicity. Suppose
@@
Consider the linear potential outcome model (7). Suppose the following exogeneity assumptions hold: $\operatorname*{cov}(\Delta X,\Delta V)=0$, $\operatorname*{cov}(\Delta W_{2},\Delta V)=0$, and $\operatorname*{cov}(\Delta W_{1},\Delta V)=0$. Then $\beta_{c}=\beta_{\text{long}}$.
@@ 4.4 Instrumental Variables
where $U$ is an unobserved scalar random variable and $(\beta_{c},\gamma_{1},\gamma_{2})$ are unknown constants. Thus $\beta_{c}$ is the causal effect of $Z$ on $Y$. In an instrumental variables analysis, this is typically called the reduced form causal effect, and $Y(z)$ are reduced form potential outcomes. Suppose $\operatorname*{cov}(Z^{\perp W_{1},W_{2}},Y(z)^{\perp W_{1},W_{2}})=0$, a linear parametric version of conditional independence between the instrument $Z$ and the unobserved $U$ given $(W_{1},W_{2})$. Then $\beta_{c}$ equals the OLS coefficient on $Z$ from the long regression of $Y$ on $(1,Z,W_{1},W_{2})$. In this model, Proposition 1 implies that $\beta_{c}$ is also obtained as the coefficient on $Z$ in the medium regression of $Y$ on $(1,Z,W_{1})$, and thus is point identified. In this case, Assumption A2 is an instrument exogeneity assumption. Our results in section 3 thus allow us to analyze the impacts of instrument exogeneity failure on conclusions about the reduced form causal effect of $Z$ on $Y$, $\beta_{c}$.
@@ 5.1 Data
Next we describe the observed covariates. We partition these covariates into $W_{1}$ and $W_{0}$ by following the implementation of Oster’s (2019) approach in Bazzi et al. (2020). $W_{1}$, the calibration covariates which are used to calibrate selection on unobservables, is a set of geographic and climate controls: Centroid Latitude, Centroid Longitude, Land area, Average rainfall, Average temperature, Elevation, Average potential agricultural yield, and Distance from the centroid to rivers, lakes, and the coast. $W_{0}$, the control covariates which are not used to calibrate selection on unobservables, includes state fixed effects. The questionnaire based outcomes use individual level data. For those analyses, we also include age, age-squared, gender, race, and survey wave fixed effects in $W_{0}$. In Bazzi et al. (2020), they were included in $W_{1}$. We instead include them in $W_{0}$ to keep the set of calibration covariates $W_{1}$ constant across the five main specifications. This allows us to directly compare the robustness of our baseline results across different specifications. We discuss the results that include state fixed effects in $W_{1}$ in Appendix B.
@@ 5.2 Baseline Model Results
[talltblr]rowsep=0pt {talltblr} [ caption = The Effect of Frontier Life on Opposition to Government Intervention and Redistribution., remarkNote = Panel A and the first row of Panel B replicate columns 1, 2, 4, 6, and 7 of table 3 in Bazzi, Fiszbein, and Gebresilasse (2020), while the second row of Panel B and Panel C are new. As in Bazzi et al. (2020), Panel B uses Oster’s rule of thumb choice $R_{Y\sim X,W_{1},W_{2}\mathrel{{\bullet}}W_{0}}^{2}=1.3\cdot\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$., ] p0.25¿ *3¿p0.15 — *2¿p0.15 Outcome: Prefers Cut Public Spending on Poor Prefers Cut Public Spending on Welfare Prefers Reduce Debt by Spending Cuts County Property Tax Rate Republican Presidential Vote Share
@@
$\widehat{\bar{r}}_{X}^{\text{ bp}}$ ($\times 100$) 2.81 3.05 5.85 72.6 80.4
$\widehat{\bar{r}}^{\text{ bp}}$ ($\times 100$) 54.1 74.4 83.9 91.8 95.9